# ORACLE: Explaining Feature Interactions in Neural Networks with ANOVA

### v2

@@ 3.5 Extracting ANOVA interaction maps and strengths
and use $\{\hat{S}_{jk}\}_{j<k}$ to rank interactions. In our experiments, these estimated ANOVA interaction maps and strengths serve as the global explanations produced by ORACLE. They are evaluated against oracle ANOVA interaction maps derived from the backbone and $P_{X}$, and compared to SHAP-family baselines in terms of interaction selection, hotspot localization, cross-backbone transfer, scale alignment, and intervention-oriented utility.
@@ 4 Theory
4.1 Oracle ANOVA interaction decomposition
We view the trained backbone $f$ as a square-integrable function on the input distribution: $X=(X_{1},\dots,X_{d})\sim P_{X}$ and $f\in L^{2}(P_{X})$. Following the functional ANOVA literature, we consider a decomposition of $L^{2}(P_{X})$ into mutually orthogonal subspaces
@@ 4.1 Oracle ANOVA interaction decomposition
where $\mathcal{H}_{0}$ is the space of constants, $\mathcal{H}_{j}$ contains main effects of $X_{j}$, $\mathcal{H}_{jk}$ contains pairwise interactions between $(X_{j},X_{k})$, and $\mathcal{R}$ collects higher-order and residual components.
Definition 4.1 (Oracle ANOVA decomposition).
For $f\in L^{2}(P_{X})$, the oracle ANOVA decomposition is
@@
Under mild regularity conditions, the ANOVA components admit conditional-expectation representations. In particular, if we define
@@
$g_{jk}(x_{j},x_{k})$ $=\mathbb{E}\bigl[f(X)\mid X_{j}=x_{j},X_{k}=x_{k}\bigr]$
$\quad-m_{j}(x_{j})-m_{k}(x_{k})-\mu.$
@@
Definition 4.2 (Oracle ANOVA interaction map and strength).
For each pair $(j,k)$, we call $g_{jk}$ the oracle ANOVA interaction map of $(X_{j},X_{k})$, and define the corresponding target ANOVA interaction strength as
@@
The decomposition in Definition 4.1 is formally identical to the classical two-factor ANOVA decomposition used in design-of-experiments: the spaces $\mathcal{H}_{j}$ and $\mathcal{H}_{jk}$ play the roles of main effects and interaction effects, respectively. The only difference is that we take $f$ to be an arbitrary learned predictor (e.g., a neural network) rather than a linear model, and we work with the data-generating distribution $P_{X}$ instead of a fixed factorial design.
4.2 ORACLE surrogate as a discrete $L^{2}$ projection
We now connect the ORACLE surrogate from Section 3 to the oracle ANOVA decomposition above. Let $q_{j}:\mathbb{R}\to\{1,\dots,L\}$ denote the binning map for feature $X_{j}$, and define the discretized variables $Z_{j}=q_{j}(X_{j})$ and $Z=(Z_{1},\dots,Z_{d})\sim P_{Z}$. On the discrete grid $\mathcal{B}_{j}=\{1,\dots,L\}$, we consider the finite-dimensional ANOVA subspaces
@@ 4.2 ORACLE surrogate as a discrete $L^{2}$ projection
spanned by contrast-coded main-effect and interaction basis functions of $Z$.
@@
Proposition 4.4 (Discrete ANOVA projection).
The minimizer $\beta^{\star}$ in Equation (1) exists and is unique under standard conditions, and the corresponding function $f^{L}$ admits a discrete ANOVA decomposition
@@
The discrete interaction components $g^{L}_{jk}$ in Proposition 4.4 can be viewed as bin-wise averages of the oracle ANOVA interaction maps $g_{jk}$ within each cell of the grid $(Z_{j},Z_{k})$. In this sense, $g^{L}_{jk}$ is the piecewise-constant $L^{2}$-projection of $g_{jk}$ onto the discretized ANOVA space.
@@
Assume the usual regularity conditions for linear regression in the ANOVA basis. Then the least-squares estimator $\hat{\beta}$ converges in probability to $\beta^{\star}$ as $n\to\infty$, and the corresponding discrete components satisfy
@@ 4.3 Consistency of interaction strengths and Top-$K$ selection
We now study the behavior of the interaction strengths induced by the discrete ANOVA maps. For a fixed grid size $L$, define the discrete interaction strength
@@
In particular, the NDCG@5 metric used in our experiments is a consistent measure of the ability of ORACLE to recover the top-$K$ target ANOVA interaction strengths.
@@ 4.4 Connection to classical $2^{5}$ factorial designs
Suppose the backbone $f$ in Equation (2) is trained to interpolate $f^{\star}$ on all $2^{5}$ design points and we apply ORACLE with $L=2$ bins per feature, using the standard contrast-coding for $\{-1,+1\}$. Then:
@@
    The population design matrix of the ANOVA surrogate coincides with the classical $2^{5}$ full factorial design matrix.
@@
    The ANOVA surrogate recovers the coefficients of $f^{\star}$ exactly, and the oracle ANOVA interaction maps satisfy
@@
Proposition 4.10 explains why ORACLE exhibits near-ideal behavior on the $2^{5}$ synthetic experiment in our evaluation: in this home-ground regime, the discrete ANOVA surrogate is algebraically identical to the classical factorial ANOVA design, so that interaction strengths, interaction maps, and cross-backbone stability metrics all align perfectly with the ground truth.
@@ Discussion.
Across the three tabular regression benchmarks, the ANOVA surrogate consistently recovered the backbone’s top interaction structure more accurately than SHAP-family interaction estimators: ORACLE attains higher NDCG@5 and Xfer-NDCG@5 in nearly all settings, together with strong Peak-IoU@0.10 and CCC, indicating better agreement with target ANOVA interaction strengths, more stable rankings across independently trained backbones, and well-localized, scale-calibrated detection of interaction hotspots. Beyond these scalar metrics, Figure 2 shows that the recovered main-effect curves match classical DoE-style factor trends, and Figure 3 in Appendix D reveals localized hotspots aligned with the detected top pairs, suggesting that the learned ANOVA structure is visually as well as numerically faithful. The intervention metric IG@K,B is more mixed: SHAP-family methods sometimes achieve similar or slightly higher IG, consistent with their game-theoretic design as marginal contribution scores tailored to removal-style interventions, whereas ORACLE is explicitly optimized for $L^{2}$ fidelity and variance decomposition and therefore yields its largest advantages on the synthetic $2^{5}$ factorial design, with still robust gains on Airfoil and kin8nm, indicating that moderately complex tabular predictors admit low-order ANOVA projections that are faithful enough for interaction benchmarking. As a complementary case study, the latent image and text experiments in Appendix E show that grid-based ANOVA surrogates are less competitive in highly entangled latent spaces, reinforcing that ORACLE is best suited to settings where features admit interpretable factorial structure.
@@
ORACLE inherits several limitations from its ANOVA and grid-based design. First, interaction maps are defined on a discretized input grid, so performance depends on binning and can suffer when many features or high resolutions lead to sparse cells; our ablation on the grid size $L$ suggests that moderate grids (e.g., $L\in\{3,5,7\}$) offer a reasonable trade-off between resolution and stability. Second, we restrict attention to pairwise interactions in regression backbones and treat the ANOVA decomposition of the trained model under the observed data distribution as the target, so higher-order effects or severe model misspecification can make these targets misaligned with domain notions of interaction. Third, SHAP baselines are evaluated under a fixed Monte Carlo budget and shared reference configuration, and alternative sampling schemes or background choices might improve their metrics and narrow the observed gaps. Finally, our main results focus on low- to medium-dimensional tabular tasks and MLP architectures, so the conclusions may not directly transfer to high-dimensional vision or language models with strong local structure or entangled latent representations.
@@
Several extensions follow naturally from these observations. Methodologically, scaling ANOVA surrogates beyond low- to medium-dimensional tabular data will likely require adaptive or sparse grids with local refinement of interaction hotspots and structured bases that better respect feature correlations, as well as extensions to classification, sequence, graph backbones, and higher-order interaction maps. On the evaluation side, combining oracle-based metrics with additional stability and robustness criteria under distribution shift, alternative training runs, or architectural changes could yield more comprehensive interaction benchmarks. Finally, closer integration with SHAP-style estimators—for example, using ANOVA maps as priors, sanity checks, or guides for targeted Monte Carlo sampling—may lead to hybrid methods that combine the structured semantics of design-of-experiments with the flexibility of Shapley-based interaction indices.
@@ A.3 Discrete ANOVA surrogate as an $L^{2}$ projection
We now formalize the population-level ANOVA surrogate obtained after discretizing the inputs and show that it coincides with an $L^{2}$ projection of the predictor onto a finite-dimensional ANOVA subspace.
@@
Recall the discretized variables $Z=(Z_{1},\dots,Z_{d})$ and the ANOVA basis
@@
where $\phi_{0}(Z)\equiv 1$, $\phi_{j}(Z)$ depends only on $Z_{j}$ and is centered, and $\phi_{jk}(Z)$ depends only on $(Z_{j},Z_{k})$ and satisfies the centering constraints in Appendix A.1. We write
$\mathcal{V}_{L}\coloneqq\big\{\beta^{\top}\phi(Z):\beta\in\mathbb{R}^{p}\big\}\subseteq L^{2}(P_{Z})$
for the corresponding finite-dimensional subspace of square-integrable functions of $Z$. Since $Z$ is a deterministic function of $X$, we can also view any element of $\mathcal{V}_{L}$ as a function of $X$ via $h(X)\coloneqq h(Z)=h\!\big(q(X)\big)$, so $\mathcal{V}_{L}$ is a closed finite-dimensional subspace of $L^{2}(P_{X})$ as well.
Assumption A.6 (Nondegenerate ANOVA design).
@@
Assumption A.6 rules out redundant basis functions and ensures that the parametrization of $\mathcal{V}_{L}$ by $\beta\in\mathbb{R}^{p}$ is well-defined.
@@
and define the population ANOVA surrogate as the minimizer of $J$.
@@
is the $L^{2}(P_{X})$-orthogonal projection of $f(X)$ onto the subspace $\mathcal{V}_{L}$, i.e.
$f_{L}=\arg\min_{g\in\mathcal{V}_{L}}\mathbb{E}\big[(f(X)-g(Z))^{2}\big],$
and the residual $f(X)-f_{L}(Z)$ is orthogonal to $\mathcal{V}_{L}$:
$\mathbb{E}\big[(f(X)-f_{L}(Z))\,g(Z)\big]=0\quad\text{for all }g\in\mathcal{V}_{L}.$
@@
Quadratic expansion of $J$ gives
@@
The function $f_{L}(Z)=\beta^{\star\top}\phi(Z)$ lies in $\mathcal{V}_{L}$ by construction. Conversely, any $g\in\mathcal{V}_{L}$ can be written as $g(Z)=\gamma^{\top}\phi(Z)$ for some $\gamma\in\mathbb{R}^{p}$, and
$\mathbb{E}\big[(f(X)-f_{L}(Z))\,g(Z)\big]=\mathbb{E}\big[(f(X)-\beta^{\star\top}\phi(Z))\,\gamma^{\top}\phi(Z)\big]=\gamma^{\top}(b-G\beta^{\star})=0,$
since $G\beta^{\star}=b$. Thus $f_{L}$ is orthogonal to $\mathcal{V}_{L}$ in the sense of the projection residual, and by the standard Hilbert-space theory of least-squares projections, $f_{L}$ is the unique $L^{2}(P_{X})$-projection of $f$ onto $\mathcal{V}_{L}$. ∎
Discrete ANOVA components.
@@
By construction of the basis, $\beta_{0}^{\star}$ is the discrete intercept, the terms $\phi_{j}(Z)^{\top}\beta_{j}^{\star}$ are centered functions of $Z_{j}$, and the terms $\phi_{jk}(Z)^{\top}\beta_{jk}^{\star}$ are centered functions of $(Z_{j},Z_{k})$ satisfying the same conditional centering constraints as $H_{j}$ and $H_{jk}$ in Appendix A.2. Hence $f_{L}$ admits a discrete ANOVA-style decomposition mirroring the oracle decomposition of Proposition A.4.
In particular, the pairwise component
@@
is the restriction of the orthogonal projection of $f$ onto the subspace spanned by $\phi_{jk}$, and can be interpreted as a binned approximation to the oracle residual surface $g_{jk}(X_{j},X_{k})$ on the $L\times L$ grid defined by $(Z_{j},Z_{k})$.

### v3

@@ 3.5 Extracting interaction maps and strengths
and use $\{\hat{S}_{jk}\}_{j<k}$ to rank interactions. In our experiments, these estimated interaction maps and strengths serve as the global explanations produced by ORACLE. They are evaluated against oracle interaction maps derived from the backbone and $P_{X}$, and compared to SHAP-family baselines in terms of interaction selection, hotspot localization, cross-backbone transfer, scale alignment, and intervention-oriented utility.
@@ 4 Theory
4.1 Oracle interaction decomposition via functional ANOVA
We view the trained backbone $f$ as a square-integrable function under the input distribution: $X=(X_{1},\dots,X_{d})\sim P_{X}$ and $f\in L^{2}(P_{X})$. Following the functional-ANOVA literature, we consider an orthogonal decomposition of $L^{2}(P_{X})$ into subspaces
@@ 4.1 Oracle interaction decomposition via functional ANOVA
where $\mathcal{H}_{0}$ is the space of constants, $\mathcal{H}_{j}$ contains main effects of $X_{j}$, $\mathcal{H}_{jk}$ contains pairwise interaction effects between $(X_{j},X_{k})$, and $\mathcal{R}$ collects higher-order and residual components.
Definition 4.1 (Oracle functional-ANOVA decomposition).
For $f\in L^{2}(P_{X})$, the oracle decomposition is
@@
Under mild regularity conditions, these components admit conditional-expectation representations. In particular, if we define
@@
$g_{jk}(x_{j},x_{k})$ $=\mathbb{E}\bigl[f(X)\mid X_{j}=x_{j},X_{k}=x_{k}\bigr]-m_{j}(x_{j})-m_{k}(x_{k})-\mu,$
@@
Definition 4.2 (Oracle interaction map and strength).
For each pair $(j,k)$, we call $g_{jk}$ the oracle interaction map of $(X_{j},X_{k})$, and define the corresponding target interaction strength as
@@
The decomposition in Definition 4.1 mirrors the classical two-factor ANOVA decomposition used in design-of-experiments: the spaces $\mathcal{H}_{j}$ and $\mathcal{H}_{jk}$ correspond to main and interaction effects, respectively. The difference is that we take $f$ to be an arbitrary learned predictor (e.g., a neural network) and we define effects with respect to the data-generating distribution $P_{X}$, rather than a fixed factorial design.
4.2 ORACLE as a discrete $L^{2}$ projection onto a factorial-effect space
We now connect the ORACLE surrogate from Section 3 to the oracle decomposition above. Let $q_{j}:\mathbb{R}\to\{1,\dots,L\}$ denote the binning map for feature $X_{j}$, and define discretized variables $Z_{j}=q_{j}(X_{j})$ and $Z=(Z_{1},\dots,Z_{d})\sim P_{Z}$. On the discrete grid $\mathcal{B}_{j}=\{1,\dots,L\}$, we consider the finite-dimensional subspaces
@@ 4.2 ORACLE as a discrete $L^{2}$ projection onto a factorial-effect space
spanned by contrast-coded main-effect and pairwise interaction-effect basis functions of $Z$.
@@
Proposition 4.4 (Discrete factorial-effect projection).
The minimizer $\beta^{\star}$ in Equation (1) exists and is unique under standard conditions, and the corresponding function $f^{L}$ admits the decomposition
@@
The discrete interaction components $g^{L}_{jk}$ in Proposition 4.4 can be interpreted as bin-wise averages of the oracle interaction maps $g_{jk}$ over each cell of the grid $(Z_{j},Z_{k})$. In this sense, $g^{L}_{jk}$ is the piecewise-constant $L^{2}$ projection of $g_{jk}$ onto the discretized factorial-effect space.
@@
Assume the usual regularity conditions for linear regression in the contrast-coded factorial basis. Then the least-squares estimator $\hat{\beta}$ converges in probability to $\beta^{\star}$ as $n\to\infty$, and the corresponding discrete components satisfy
@@ 4.3 Consistency of interaction strengths and Top-$K$ selection
We now study the behavior of interaction strengths induced by the discrete interaction maps. For a fixed grid size $L$, define the discrete interaction strength
@@
In particular, the NDCG@5 metric used in our experiments is a consistent measure of the ability of ORACLE to recover the top-$K$ target interaction strengths.
@@ 4.4 Connection to classical $2^{5}$ factorial designs
Suppose the backbone $f$ in Equation (2) is trained to interpolate $f^{\star}$ on all $2^{5}$ design points and we apply ORACLE with $L=2$ bins per feature, using the standard contrast coding for $\{-1,+1\}$. Then:
@@
    The population design matrix of the factorial-effect surrogate coincides with the classical $2^{5}$ full factorial design matrix.
@@
    The factorial-effect surrogate recovers the coefficients of $f^{\star}$ exactly, and the oracle interaction maps satisfy
@@
Proposition 4.10 explains why ORACLE exhibits near-ideal behavior on the $2^{5}$ synthetic experiment in our evaluation: in this home-ground regime, the discrete ANOVA surrogate (i.e., the grid-based orthogonal factorial expansion) is algebraically identical to the classical $2^{5}$ factorial ANOVA design. Consequently, the induced interaction strengths, interaction maps, and cross-backbone stability metrics align with the ground truth.
@@ Discussion.
Across the three tabular regression benchmarks, the factorial surrogate consistently recovers the backbone’s top interaction structure more accurately than SHAP-family interaction estimators: ORACLE attains higher NDCG@5 and Xfer-NDCG@5 in nearly all settings, together with strong Peak-IoU@0.10 and CCC. These results indicate better agreement with the target interaction strengths, more stable interaction rankings across independently trained backbones, and well-localized, scale-calibrated detection of interaction hotspots. Beyond these scalar metrics, Figure 2 shows that the recovered main-effect curves match classical DoE-style factor trends, and Figure 3 (Appendix D) reveals localized hotspots aligned with the detected top pairs, suggesting that the recovered low-order structure is visually as well as numerically faithful. The intervention metric IG@K,B is more mixed: SHAP-family methods sometimes achieve similar or slightly higher IG, consistent with their game-theoretic design as marginal-contribution scores tailored to removal-style interventions. In contrast, ORACLE is optimized for $L^{2}$ fidelity of effect surfaces and for capturing low-order structure on the grid, and therefore shows its largest advantages on the synthetic $2^{5}$ factorial design, with still robust gains on Airfoil and kin8nm. Taken together, these findings suggest that moderately complex tabular predictors often admit low-order projections that are faithful enough for interaction benchmarking. As a complementary case study, the latent image and text experiments in Appendix E show that grid-based factorial surrogates are less competitive in highly entangled latent spaces, reinforcing that ORACLE is best suited to settings where features admit interpretable factorial structure.
@@
ORACLE inherits several limitations from its grid-based factorial design. First, interaction maps are defined on a discretized input grid, so performance depends on binning and can degrade when many features or high resolutions lead to sparse cells; our ablation on the grid size $L$ suggests that moderate grids (e.g., $L\in\{3,5,7\}$) offer a reasonable trade-off between resolution and stability. Second, we restrict attention to pairwise interactions in regression backbones and treat the functional-ANOVA decomposition of the trained model under the observed data distribution as the target; higher-order effects, strong feature dependence, or model misspecification can therefore make these targets misaligned with domain notions of interaction. Third, SHAP baselines are evaluated under a fixed Monte Carlo budget and shared reference configuration, and alternative sampling schemes or background choices might improve their metrics and narrow the observed gaps. Finally, our main results focus on low- to medium-dimensional tabular tasks and MLP architectures, so the conclusions may not directly transfer to high-dimensional vision or language models with strong local structure or entangled latent representations.
@@
Several extensions follow naturally from these observations. Methodologically, scaling factorial surrogates beyond low- to medium-dimensional tabular data will likely require adaptive or sparse grids with local refinement around interaction hotspots and structured bases that better respect feature dependence, as well as extensions to classification, sequence, and graph backbones and to higher-order interaction maps. On the evaluation side, combining oracle-based metrics with additional stability and robustness criteria under distribution shift, alternative training runs, or architectural changes could yield more comprehensive interaction benchmarks. Finally, closer integration with SHAP-style estimators—for example, using interaction maps as priors, sanity checks, or guides for targeted Monte Carlo sampling—may lead to hybrid methods that combine the structured semantics of design-of-experiments with the flexibility of Shapley-based interaction indices.
@@ A.3 Discrete ANOVA surrogate as an $L^{2}$ projection
We now formalize the population-level surrogate obtained after discretizing the inputs and show that it coincides with the $L^{2}(P_{X})$ orthogonal projection of the predictor onto a finite-dimensional ANOVA (factorial-effect) subspace.
@@
Recall the discretized variables $Z=(Z_{1},\dots,Z_{d})$ and the contrast-coded basis
@@
where $\phi_{0}(Z)\equiv 1$, each $\phi_{j}(Z)$ depends only on $Z_{j}$ and is centered, and each $\phi_{jk}(Z)$ depends only on $(Z_{j},Z_{k})$ and satisfies the same centering constraints as in Appendix A.1. Define the finite-dimensional subspace of square-integrable functions of $Z$ by
$\mathcal{V}_{L}^{(Z)}\coloneqq\big\{\beta^{\top}\phi(Z):\beta\in\mathbb{R}^{p}\big\}\subseteq L^{2}(P_{Z}).$
Since $Z=q(X)$ is a deterministic function of $X$, we also consider the induced (substituted) subspace in $L^{2}(P_{X})$,
$\mathcal{V}_{L}\coloneqq\big\{g\!\big(q(X)\big):g\in\mathcal{V}_{L}^{(Z)}\big\}=\big\{\beta^{\top}\phi\!\big(q(X)\big):\beta\in\mathbb{R}^{p}\big\}\subseteq L^{2}(P_{X}).$
Because $\mathcal{V}_{L}$ is finite-dimensional, it is closed in $L^{2}(P_{X})$.
Assumption A.6 (Nondegenerate factorial-effect design).
@@
Assumption A.6 rules out redundant basis functions and ensures that the parametrization of $\mathcal{V}_{L}^{(Z)}$ by $\beta\in\mathbb{R}^{p}$ is well-defined.
@@
and define the population surrogate as the minimizer of $J$.
@@
induces an element $f_{L}\!\big(q(X)\big)\in\mathcal{V}_{L}$ and equals the $L^{2}(P_{X})$-orthogonal projection of $f(X)$ onto $\mathcal{V}_{L}$, i.e.
$f_{L}\!\big(q(X)\big)=\arg\min_{h\in\mathcal{V}_{L}}\mathbb{E}\big[(f(X)-h(X))^{2}\big].$
Equivalently, the residual is orthogonal to $\mathcal{V}_{L}$:
$\mathbb{E}\big[(f(X)-f_{L}(Z))\,g(Z)\big]=0\quad\text{for all }g\in\mathcal{V}_{L}^{(Z)}.$
@@
A quadratic expansion of $J$ gives
@@
Let $f_{L}(Z)=\beta^{\star\top}\phi(Z)\in\mathcal{V}_{L}^{(Z)}$. For any $g\in\mathcal{V}_{L}^{(Z)}$, write $g(Z)=\gamma^{\top}\phi(Z)$ for some $\gamma\in\mathbb{R}^{p}$. Then
$\mathbb{E}\big[(f(X)-f_{L}(Z))\,g(Z)\big]=\gamma^{\top}\big(b-G\beta^{\star}\big)=0,$
since $G\beta^{\star}=b$. Thus the residual is orthogonal to $\mathcal{V}_{L}^{(Z)}$, and therefore $f_{L}\!\big(q(X)\big)$ is the unique $L^{2}(P_{X})$ projection of $f$ onto $\mathcal{V}_{L}$. ∎
Discrete effect components.
@@
By construction of the basis, $\beta_{0}^{\star}$ is the discrete intercept, each $\phi_{j}(Z)^{\top}\beta_{j}^{\star}$ is a centered function of $Z_{j}$, and each $\phi_{jk}(Z)^{\top}\beta_{jk}^{\star}$ is a centered function of $(Z_{j},Z_{k})$ satisfying the same conditional-centering constraints as in Appendix A.2. Hence $f_{L}$ admits a discrete decomposition that mirrors the oracle functional-ANOVA decomposition.
In particular, the pairwise term
@@
is the interaction table associated with the $(j,k)$ block of the projected surrogate $f_{L}$. It can be interpreted as a binned approximation to the oracle interaction map $g_{jk}(X_{j},X_{k})$ on the $L\times L$ grid defined by $(Z_{j},Z_{k})$.